OpenStudy (anonymous):

Four coins are tossed simultaneously. What is the probability that all 4 coins will land on heads? Type your answer as a percent

OpenStudy (anonymous):

50%

OpenStudy (anonymous):

That is incorrect

OpenStudy (anonymous):

@LegendarySadist How so?

OpenStudy (anonymous):

what do I do idk I thought it was close to that

OpenStudy (anonymous):

You multiply the percentages together. What's the chance of getting a head on a coin flip?

OpenStudy (anonymous):

50% @LegendarySadist

OpenStudy (anonymous):

So they all have a 50% chance of being heads. Now multiply the percentages together. It might be easier to write it as \(\large \frac{1}{2}\) when you multiply them.

OpenStudy (anonymous):

50% is wrong snuffitaly I checked it on there

Vocaloid (vocaloid):

the probability of ONE coin being heads is 50% however, we have FOUR coins here, so we must multiply 50% 4 times (0.5)*(0.5)*(0.5)*(0.5) does that make a little more sense?

OpenStudy (anonymous):

Ok thanks @LegendarySadist and @Vocaloid

OpenStudy (anonymous):

So would it be 200?

Vocaloid (vocaloid):

um, not quite... put in (0.5)*(0.5)*(0.5)*(0.5) into your calculator and tell me what you get

Vocaloid (vocaloid):

* means multiply

OpenStudy (anonymous):

\[\large \frac{1}{2} \times \frac{1}{2} \times \frac{1}{2} \times \frac{1}{2}\]

OpenStudy (anonymous):

0.0625 is what i got

Vocaloid (vocaloid):

right, so convert that decimal to a percent by multiplying by 100

OpenStudy (anonymous):

6.25

Vocaloid (vocaloid):

right, so our final answer is 6.25%

OpenStudy (anonymous):

ok thanks